# CaRiNG: Learning Temporal Causal Representation under Non-Invertible Generation Process

### v1

@@ A.4 Identifiability Benefits from Non-Stationarity
Drawing upon the arguments in the proof of Theorem 1, given that the components of $\hat{\mathbf{z}}_{t}$ are mutually independent conditional on $\hat{\mathbf{z}}_{t-1}$, we know that for $i\neq j$,
$\frac{\partial^{2}\log p(\hat{\mathbf{z}}_{t}\,|\,\hat{\mathbf{z}}_{t-1};\mathbf{u})}{\partial\hat{z}_{it}\partial\hat{z}_{jt}}=\sum_{k=1}^{n}\Big{(}\frac{\partial^{2}\eta_{kt}(\mathbf{u})}{\partial z_{kt}^{2}}\cdot\mathbf{H}_{kit}\mathbf{H}_{kjt}+\frac{\partial\eta_{kt}(\mathbf{u})}{\partial z_{kt}}\cdot\frac{\partial\mathbf{H}_{kit}}{\partial\hat{z}_{jt}}\Big{)}-\frac{\partial^{2}\log|\mathbf{H}_{t}|}{\partial\hat{z}_{it}\partial\hat{z}_{jt}}\equiv 0.$ (28)
In contrast to Eq 22, we now allow $p(\hat{\mathbf{z}}_{t}\,|\,\hat{\mathbf{z}}_{t-1})$ to depend on $\mathbf{u}$. Given that the aforementioned equation is always 0, its partial derivative w.r.t. $z_{l,t-1}$ yields
$\frac{\partial^{3}\log p(\hat{\mathbf{z}}_{t}\,|\,\hat{\mathbf{z}}_{t-1};\mathbf{u})}{\partial\hat{z}_{it}\partial\hat{z}_{jt}\partial z_{l,t-1}}$ $=\sum_{k=1}^{n}\Big{(}\frac{\partial^{3}\eta_{kt}(\mathbf{u})}{\partial z_{kt}^{2}\partial z_{l,t-1}}\cdot\mathbf{H}_{kit}\mathbf{H}_{kjt}+\frac{\partial^{2}\eta_{kt}(\mathbf{u})}{\partial z_{kt}\partial z_{l,t-1}}\cdot\frac{\partial\mathbf{H}_{kit}}{\partial\hat{z}_{jt}}\Big{)}\equiv 0.$ (29)
Similarly, when using varied values for $\mathbf{u}$ in Eq 28, computing the difference between these instances yields
$\frac{\partial^{2}\log p(\hat{\mathbf{z}}_{t}\,|\,\hat{\mathbf{z}}_{t-1};{u}_{r+1})}{\partial\hat{z}_{it}\partial\hat{z}_{jt}}-\frac{\partial^{2}\log p(\hat{\mathbf{z}}_{t}\,|\,\hat{\mathbf{z}}_{t-1};{u}_{r+1})}{\partial\hat{z}_{it}\partial\hat{z}_{jt}}$
$=$ $\sum_{k=1}^{n}\Big{[}\Big{(}\frac{\partial^{2}\eta_{kt}({u}_{r+1})}{\partial z_{kt}^{2}}-\frac{\partial^{2}\eta_{kt}({u}_{r})}{\partial z_{kt}^{2}}\Big{)}\cdot\mathbf{H}_{kit}\mathbf{H}_{kjt}+\Big{(}\frac{\partial\eta_{kt}(u_{r+1})}{\partial z_{kt}}-\frac{\partial\eta_{kt}(u_{r})}{\partial z_{kt}}\Big{)}\cdot\frac{\partial\mathbf{H}_{kit}}{\partial\hat{z}_{jt}}\Big{]}\equiv 0.$ (30)
Therefore, if $\mathbf{s}_{kt}$ and $\mathring{\mathbf{s}}_{kt}$, for $k=1,2,...,n$, are linearly independent, $\mathbf{H}_{kit}\mathbf{H}_{kjt}$ has to be zero for all $k$ and $i\neq j$. Building on the insights from the proof of Theorem 1, $\hat{\mathbf{z}}_{t}$ is compelled to be a permutation of a component-wise invertible transformation of $\mathbf{z}_{t}$. ∎
Appendix B Experiment Settings
B.1 Reproducibility
All experiments are done in a GPU workstation with CPU: Intel(R) Xeon(R) Platinum 8168 CPU @ 2.70GHz, GPU: Tesla V100. The source code and the generated data for the simulation experiments are attached in the supplementary materials.
B.2 Synthetic Dataset Generation
@@
For NG, we set the transition lag as $\tau=1$. We first generate $10,000$ data points from uniform distribution as the initial state $\mathbf{z}_{0}\sim U(0,1)$. For $t=1,\cdots,9$, each latent variable $\mathbf{z}_{t}$ will be generated from the proceeding latent variable $\mathbf{z}_{t-1}$ through a nonlinear function $\mathbf{f}$ with a non-additive zero-biased Gaussian noise $\mathbf{\epsilon}_{t}$ ($\sigma=0.1$), i.e., $\mathbf{z}_{t}=\mathbf{f}(\mathbf{z}_{t},\epsilon_{t})$. To introduce the non-invertibility, the mixing function $\mathbf{g}$ leverages only the first two entries of the latent variables to generate the 2-d observation $\mathbf{z}_{t}=\mathbf{g}(x_{1,t},x_{2,t})\in\mathcal{Z}\subseteq\mathbb{R}^{2}$.
@@
For UG-TDMP, we set the transition lag as $\tau=1$ and mixing lag $r=2$. Similar to the Non-invertible Generation scenario, we generate the initial states from uniform distribution and the subsequent latent variables following a nonlinear transition function. The noise is also introduced in a nonlinear Gaussian ($\sigma=0.1$) way. The mixing process is a nonlinear function with regard to $\mathbf{z}_{t}$ plus a side information from previous steps $\mathbf{z}_{t-1:t-2}$, i.e.,
$\mathbf{x}_{t}=A_{3\times 3}\cdot\sigma\big{(}B_{3\times 3}\cdot\sigma(C_{3\times 3}\cdot\mathbf{z}_{t})\big{)}+\begin{bmatrix}0\\ 0\\ D_{3\times 1}\mathbf{z}_{t-1}+E_{3\times 1}\mathbf{z}_{t-2}\end{bmatrix},$ (31)
where $\sigma$ refers to the ReLU function and the capital characters refer to matrices. Note that we make two modifications to show the advantage of CaRiNG. The reason we consider larger mixing lag is that it is a much more difficult scenario to handle, with more distribution from the mixing process and less dynamic information from transition. We run experiments in both scenarios with different transition and mixing lag. Besides, we also find out that even without time-lagged latent variables in the decoder, it leads to a smaller model that is more stable and easy to train. Refer to Table 3 for a detailed ablation study.
@@
Table 3: Ablation study on different settings for UG-TDMP. (a) The second column is a more difficult scenario compared to the first, where the performance of CaRiNG remains good while that of baseline decreases significantly. (b) Omit the time-lagged latent variables in the decoder will not damage the performance much, but one can enjoy the benefits from a much simpler model.
@@
B.3 Implementation Details
B.3.1 Synthetic Data
@@
To implement the Sequence-to-Step encoder, we leverage the torch.unfold to generate the nesting observations. Let us denote $\mathbf{x}_{t}^{(\mu)}=[\mathbf{x}_{t},\cdots,\mathbf{x}_{t-\mu}]$ as inputs. For the time steps that do not exist, we simply pad them with zero. Refer to Table 4 for detailed network architecture.
@@
Table 4: Architecture details. BS: batch size, T: length of time series, i_dim: input dimension, o_dim: output dimension, z_dim: latent dimension, LeakyReLU: Leaky Rectified Linear Unit.
@@ 4. Modular Prior Nonlinear Transition Prior Network
B.3.2 Real-world Dataset
Network Architecture
We choose HCRN (Le et al., 2020) (without classification head) as the encoder backbone of CaRiNG on the real-world dataset: SUTD-TrafficQA. Given that HCRN is an encoder that calculates the cross attention between visual input and text input sequentially, we apply a decoder, which shares the same structure as the Step-to-Step Decoder shown in Table 4 to reconstruct the visual feature embedded with the temporal information. As it goes to transition prior, we use the Modular Prior shown in Table 4. This encoder-decoder structure can guide the model to learn the hidden representation with identifiable guarantees under the non-invertible generation process.
Appendix C More visualization results on real-world data
As shown Figure 5, we provide some positive examples and also fail cases to analyze our model. From the top two examples, we can find that our method can solve the occlusions well. From the bottom right one, we find that our model can solve the blurred situation. However, when the alignment between visual and textual domains is difficult. The model may fail.
Figure 5: Qualitative resutls on SUTD-TrafficQA dataset. We provide some positive examples and also fail cases to analyze our model.
Appendix D More experimental results
D.1 Computation Cost Comparison
We provide the comparisons between the computational cost of the CaRiNG model compared to HCRN to analyze our efficiency. As shown in Table 5, we provide a detailed comparison of the number of parameters, training time, and inference efficiency. It is important to note that while the CaRiNG model requires a longer training time due to the application of normalizing flow for calculating the Jacobian matrix, its inference efficiency remains on par with HCRN, as the normalizing flow is utilized only for calculating KL loss and not during inference.
@@ D.1 Computation Cost Comparison
Table 5: Comparative Analysis of HCRN and CaRiNG Models
@@
D.2 Evaluation of Identifiability in the QA Benchmark
In the context of real-world applications, particularly in scenarios lacking ground truth for rigorous metrics like MCC, alternative evaluation strategies become essential. we leverage proxy metrics to assess the performance of the proposed algorithm, focusing on two pivotal aspects: disentanglement and reconstruction ability of the learned representations. Intuitively, as delineated in Theorem 2 and detailed in Section 4, a representation can be considered identifiable if it possesses the dual capability of fully reconstructing the observation while also achieving disentanglement. Thus, as a supplement to the accuracy we used before, we benchmark disentanglement and reconstruction ability as side evidence to support that the improvement is caused by better identifiability.
We use the ELBO loss as a proxy metric to evaluate the identifiability. Figure 6 illustrates our method’s performance compared to the baseline TDRL method. The results clearly show that our approach exhibits superior disentanglement and reconstruction abilities. This evidence suggests that the advantage of our proposed algorithm is primarily attributed to its enhanced identifiability and effective disentanglement of data representations.
Figure 6: Comparative analysis of disentanglement and reconstruction abilities of different methods.
D.3 Parameter analysis on $\tau$ in SUTD-TrafficQA
@@
Table 6: Parameter analysis results of $\tau$ on model accuracy in the SUTD-TrafficQA dataset.
D.4 Model Selection with Varying $\mu$
In this subsection, we discuss a preliminary experiment that was instrumental in the model selection process for our application in the NG-TDMP settings. The experiment focused on evaluating the performance of the model with varying lengths of time lag $\mu$.
$\mu$ 3 4 5
Accuracy (%) 0.88 0.92 0.92
Table 7: Impact of varying $\mu$ on model performance in NG-TDMP settings.
Appendix E Related Work
E.1 Causal Discovery with Latent Variables
@@
E.2 Nonlinear ICA for Time Series Data
Nonlinear ICA represents an alternative methodology to identify latent causal variables within time series data. Such methods leverage auxiliary data—like class labels and domain indices—and impose independence constraints to facilitate the identifiability of latent variables. To illustrate: Time-contrastive learning (TCL (Hyvarinen & Morioka, 2016)) adopts the independent sources premise and capitalizes on the variability in variance across different data segments. Furthermore, Permutation-based contrastive (PCL (Hyvarinen & Morioka, 2017)) puts forth a learning paradigm that distinguishes genuine independent sources from their permuted counterparts. Furthermore, i-VAE (Khemakhem et al., 2020) utilizes deep neural networks, VAEs, to closely approximate the joint distribution encompassing observed and auxiliary non-stationary regimes. Recent work, exemplified by LEAP (Yao et al., 2022b), has tackled both stationary and non-stationary scenarios in tandem. In the stationary context, LEAP postulates a linear non-Gaussian generative process. For the non-stationary context, it assumes a nonlinear generative process, gaining leverage from auxiliary variables. Advancing beyond LEAP, TDRL (Yao et al., 2022a) initially extends the linear non-Gaussian generative assumption to a nonlinear formulation for stationary scenarios. Subsequently, it broadens the non-stationary framework to accommodate structural shifts, global alterations, and combinations thereof. Additionally, CITRIS (Lippe et al., 2022b, a) champions the use of intervention target data to precisely identify scalar and multi-dimensional latent causal factors. However, a common thread across these methodologies is the presumption of an invertible generative process, a stance that often deviates from the realities of actual data.
E.3 Temporal modeling
@@
Appendix F Broader Impacts, Limitation, and Future Work

### v2

@@ Acknowledgments
For any $t$ we have
$\mathbf{z}_{t}$ $=\mathbf{m}(\mathbf{x}_{t:t-\mu})$ (A24)
as well as $\mathbf{\hat{z}}_{t}=\mathbf{\hat{m}}(\mathbf{g}(\mathbf{z}_{t},\mathbf{m}(\mathbf{x}_{t-1:t-\mu-1}),\cdots,\mathbf{m}(\mathbf{x}_{t-r:t-\mu-r})),\mathbf{x}_{t-1:t-\mu})$ similarly. Thus, we have an unified partially invertible function $\mathbf{z}_{t}=\mathbf{h}(\mathbf{\hat{z}}_{t}|\mathbf{x}_{t-1:t-\mu-r})$ where $\mathbf{h}=\mathbf{m}\circ\hat{\mathbf{g}}$ with Jacobian $\frac{\partial\mathbf{z}_{t}}{\partial\mathbf{\hat{z}}_{t}}=\mathbf{H}_{t}(\mathbf{\hat{z}}_{t};\mathbf{x}_{t-1:t-\mu-r})$. Let us consider the mapping from joint distribution $(\mathbf{\hat{z}}_{t},\mathbf{x}_{t-1:t-\mu-r-1})$ to $(\mathbf{z}_{t},\mathbf{x}_{t-1:t-\mu-r-1})$, i.e.,
$P(\mathbf{z}_{t},\mathbf{x}_{t-1:t-\mu-r-1})=P(\mathbf{\hat{z}}_{t},\mathbf{x}_{t-1:t-\mu-r-1})\,/\,|\mathbf{J}_{t}|,$ (A25)
where
$\mathbf{J}_{t}=\begin{bmatrix}\frac{\partial\mathbf{z}_{t}}{\partial\mathbf{\hat{z}}_{t}}&\mathbf{0}\\ *&\mathbf{I}\end{bmatrix},$ (A26)
which is a lower triangle matrix, where $\mathbf{I}$ infers eye matrix and * infers any possible matrix. Thus, we have determinant $|\mathbf{J}_{t}|=|\frac{\partial\mathbf{z}_{t}}{\partial\mathbf{\hat{z}}_{t}}|=|\mathbf{H}_{t}|$. Dividing both sides of Eq A25 by $P(\mathbf{x}_{t-1:t-\mu-r-1},u_{r})$ gives
$\textbf{LHS}=P(\mathbf{z}_{t}|\mathbf{x}_{t-1:t-\mu-r-1},u_{r})=P(\mathbf{z}_{t}|\mathbf{z}_{t-1:t-r-1},u_{r}),$ (A27)
since $\mathbf{z}_{t}$ and $\mathbf{x}_{t-1:t-\mu-r-1}$ are independent conditioned on $\mathbf{z}_{t-1:t-r-1}$ with any auxiliary variable $u_{r}$. Similarly, $\textbf{RHS}=P(\mathbf{\hat{z}}_{t}|\mathbf{x}_{t-1:t-\mu-r-1},u_{r})=P(\mathbf{\hat{z}}_{t}|\mathbf{\hat{z}}_{t-r-1},u_{r})$ holds true as well, which yields to
$P(\mathbf{z}_{t}|\mathbf{z}_{t-1:t-r-1},u_{r})=P(\mathbf{\hat{z}}_{t}|\mathbf{\hat{z}}_{t-1:t-r-1},u_{r})\,/\,|\mathbf{H}_{t}|.$ (A28)
With conditional independence, we have
$\frac{\partial^{2}\log p(\hat{\mathbf{z}}_{t}\,|\,\hat{\mathbf{z}}_{t-1:t-r-1},u_{r})}{\partial\hat{z}_{it}\partial\hat{z}_{jt}}=0.$ (A29)
Referencing Eq A28, it gets expressed as:
$\log p(\hat{\mathbf{z}}_{t}\,|\,\hat{\mathbf{z}}_{t-1:t-r-1},u_{r})=\log p({\mathbf{z}}_{t}\,|\,{\mathbf{z}}_{t-1:t-r-1},u_{r})+\log|\mathbf{H}_{t}|=\sum_{k=1}^{n}\eta_{kt}(u_{r})+\log|\mathbf{H}_{t}|.$ (A30)
The second-order derivative is
$\sum_{k=1}^{n}\Big{(}\frac{\partial^{2}\eta_{kt}(u_{r})}{\partial z_{kt}^{2}}\cdot\mathbf{H}_{kit}\mathbf{H}_{kjt}+\frac{\partial\eta_{kt}(u_{r})}{\partial z_{kt}}\cdot\frac{\partial\mathbf{H}_{kit}}{\partial\hat{z}_{jt}}\Big{)}+\frac{\partial^{2}\log|\mathbf{H}_{t}|}{\partial\hat{z}_{it}\partial\hat{z}_{jt}}\equiv 0.$ (A31)
The right-hand side of the presented equation consistently equals 0. Therefore, for each index $l$ ranging from 1 to $n$, and every associated value of $z_{l,t-r-1}$, its partial derivative with respect to $z_{l,t-r-1}$ remains 0. That is,
$\sum_{k=1}^{n}\Big{(}\frac{\partial^{3}\eta_{kt}(u_{r})}{\partial z_{kt}^{2}\partial z_{l,t-r-1}}\cdot\mathbf{H}_{kit}\mathbf{H}_{kjt}+\frac{\partial^{2}\eta_{kt}(u_{r})}{\partial z_{kt}\partial z_{l,t-r-1}}\cdot\frac{\partial\mathbf{H}_{kit}}{\partial\hat{z}_{jt}}\Big{)}\equiv 0,$ (A32)
where we leveraged the fact that entries of $\mathbf{H}_{t}$ do not depend on $z_{l,t-r-1}$.
Again start from Eq A31. Using the fact that $\mathbf{H}_{t}$ is not affected by the auxiliary variable, we can subtract the equation with $u_{0}$ from that of $u_{r}$. We have
$0=\sum_{k=1}^{n}\Big{(}\Big{(}\frac{\partial^{2}\eta_{kt}(u_{r})}{\partial z_{kt}^{2}}-\frac{\partial^{2}\eta_{kt}(u_{0})}{\partial z_{kt}^{2}}\Big{)}\cdot\mathbf{H}_{kit}\mathbf{H}_{kjt}+\Big{(}\frac{\partial\eta_{kt}(u_{r})}{\partial z_{kt}}-\frac{\partial\eta_{kt}(u_{0})}{\partial z_{kt}}\Big{)}\cdot\frac{\partial\mathbf{H}_{kit}}{\partial\hat{z}_{jt}}\Big{)}.$ (A33)
Considering any given value of $\mathbf{z}_{t}$, there exists at least $2n$ different values of $\mathbf{s}_{lt}$ or $\mathring{\mathbf{s}}_{t}$, which corresponds to Eq A32 and Eq A33 respectively, such that they are linearly independent. To make the above equation hold true, one has to set $\mathbf{H}_{kit}\mathbf{H}_{kjt}=0$ or $i\neq j$. In other words, each row of $\mathbf{H}_{t}$ consists of at most a single non-zero entry, and $\mathbf{z}_{t}$ must be a component-wise transformation of a permuted version of $\mathbf{\hat{z}}_{t}$. ∎
Appendix A2 Synthetic experiments
A2.1 Synthetic Dataset Generation
@@
For NG, we set the transition lag as $\tau=1$. We first generate $10,000$ data points from a uniform distribution as the initial state $\mathbf{z}_{0}\sim U(0,1)$. For $t=1,\cdots,9$, each latent variable $\mathbf{z}_{t}$ will be generated from the proceeding latent variable $\mathbf{z}_{t-1}$ through a nonlinear function $\mathbf{f}$ with a non-additive zero-biased Gaussian noise $\mathbf{\epsilon}_{t}$ ($\sigma=0.1$), i.e., $\mathbf{z}_{t}=\mathbf{f}(\mathbf{z}_{t},\epsilon_{t})$. To introduce the non-invertibility, the mixing function $\mathbf{g}$ leverages only the first two entries of the latent variables to generate the 2-d observation $\mathbf{z}_{t}=\mathbf{g}(x_{1,t},x_{2,t})\in\mathcal{Z}\subseteq\mathbb{R}^{2}$.
@@
For UG-TDMP, we set the transition lag as $\tau=1$ and mixing lag $r=2$. Similar to the Non-invertible Generation scenario, we generate the initial states from a uniform distribution and the subsequent latent variables following a nonlinear transition function. The noise is also introduced in a nonlinear Gaussian ($\sigma=0.1$) way. The mixing process is a nonlinear function with regard to $\mathbf{z}_{t}$ plus a side information from previous steps $\mathbf{z}_{t-1:t-2}$, i.e.,
$\mathbf{x}_{t}=A_{3\times 3}\cdot\sigma\big{(}B_{3\times 3}\cdot\sigma(C_{3\times 3}\cdot\mathbf{z}_{t})\big{)}+\begin{bmatrix}0\\ 0\\ D_{3\times 1}\mathbf{z}_{t-1}+E_{3\times 1}\mathbf{z}_{t-2}\end{bmatrix},$ (A34)
where $\sigma$ refers to the ReLU function and the capital characters refer to matrices. Note that we make two modifications to show the advantage of CaRiNG. The reason we consider larger mixing lag is that it is a much more difficult scenario to handle, with more distribution from the mixing process and less dynamic information from transition. We run experiments in both scenarios with different transition and mixing lag. Besides, we also find out that even without time-lagged latent variables in the decoder, it leads to a smaller model that is more stable and easy to train. Refer to Table A1 for a detailed ablation study.
@@
Table A1: Ablation study on different settings for UG-TDMP. (a) The second column is a more difficult scenario compared to the first, where the performance of CaRiNG remains good while that of baseline decreases significantly. (b) Omit the time-lagged latent variables in the decoder will not damage the performance much, but one can enjoy the benefits from a much simpler model.
@@
A2.2 Implementation Details
@@
To implement the Sequence-to-Step encoder, we leverage the torch.unfold to generate the nesting observations. Let us denote $\mathbf{x}_{t}^{(\mu)}=[\mathbf{x}_{t},\cdots,\mathbf{x}_{t-\mu}]$ as inputs. For the time steps that do not exist, we simply pad them with zero. Refer to Table A2 for detailed network architecture.
@@
Table A2: Architecture details. BS: batch size, T: length of time series, i_dim: input dimension, o_dim: output dimension, z_dim: latent dimension, LeakyReLU: Leaky Rectified Linear Unit.
@@ 4. Modular Prior Nonlinear Transition Prior Network
Table A3: MCC scores of synthetic datasets with higher dimension.
Dimension CaRiNG TDRL
6 0.9199 0.6329
12 0.9366 0.6155
18 0.7175 0.5265
A2.3 Exploration on higher dimension.
To demonstrate the scalability of our method, we have included experiments with higher dimensions. We keep the experimental setup consistent with NG and set the dimensions of latent variables to be $6,12,18$ and that of observation to be $4,8,12$, respectively. The transition function is a permutation function with a shift of $2,4,6$ dimensions respectively. As shown in Table A3, CaRiNG can achieve a consistent improvement over the baseline TDRL when using various dimensions. When the dimension is too high, although the performance of both CaRiNG and TDRL drops because of the complexity of the model, we demonstrate that CaRiNG still benefits from contextual information. This indicates that CaRiNG is scalable and robust to the dimensionality of the latent variables.
A2.4 Model Selection with Varying $\mu$
In this subsection, we will discuss the preliminary experiment that was instrumental in the model selection process for our application in the NG-TDMP settings. The experiment focused on evaluating the performance of the model with varying lengths of time lag $\mu$.
$\mu$ 3 4 5
Accuracy (%) 0.88 0.92 0.92
Table A4: Impact of varying $\mu$ on model performance in NG-TDMP settings.
Appendix A3 Real-world Experiments on TrafficQA
A3.1 Implementation Details
We choose HCRN (Le et al., 2020) (without classification head) as the encoder backbone of CaRiNG on the real-world dataset: SUTD-TrafficQA. Given that HCRN is an encoder that calculates the cross attention between visual input and text input sequentially, we apply a decoder, which shares the same structure as the Step-to-Step Decoder shown in Table A2 to reconstruct the visual feature embedded with the temporal information. As it goes to transition prior, we use the Modular Prior shown in Table A2. This encoder-decoder structure can guide the model to learn the hidden representation with identifiable guarantees under the non-invertible generation process.
A3.2 More Qualitative Results
As shown in Figure A1, we provide some positive examples and also fail cases to analyze our model. From the top two examples, we can find that our method can solve the occlusions well. From the bottom right one, we find that our model can solve the blurred situation. However, when the alignment between visual and textual domains is difficult. The model may fail.
Figure A1: Qualitative results on SUTD-TrafficQA dataset. We provide some positive examples and also fail cases to analyze our model.
A3.3 Computation Cost Comparison
We provide the comparisons between the computational cost of the CaRiNG model and HCRN to analyze our efficiency. As shown in Table A5, we provide a detailed comparison of the number of parameters, training time, and inference efficiency. It is important to note that while the CaRiNG model requires a longer training time due to the application of normalizing flow for calculating the Jacobian matrix, its inference efficiency remains on par with HCRN, as the normalizing flow is utilized only for calculating KL loss and not during inference.
@@
Table A5: Comparative Analysis of HCRN and CaRiNG Models
@@
A3.4 Evaluation of Identifiability in the QA Benchmark
In the context of real-world applications, particularly in scenarios lacking ground truth for rigorous metrics like MCC, alternative evaluation strategies become essential. we leverage proxy metrics to assess the performance of the proposed algorithm, focusing on two pivotal aspects: disentanglement and reconstruction ability of the learned representations. Intuitively, as delineated in Theorem A1 and detailed in Section 4, a representation can be considered identifiable if it possesses the dual capability of fully reconstructing the observation while also achieving disentanglement. Thus, as a supplement to the accuracy we used before, we benchmark disentanglement and reconstruction ability as side evidence to support that the improvement is caused by better identifiability.
We use the ELBO loss as a proxy metric to evaluate the identifiability. Figure A2 illustrates our method’s performance compared to the baseline TDRL method. The results clearly show that our approach exhibits superior disentanglement and reconstruction abilities. This evidence suggests that the advantage of our proposed algorithm is primarily attributed to its enhanced identifiability and effective disentanglement of data representations.
Figure A2: Comparative analysis of disentanglement and reconstruction abilities of different methods.
A3.5 Parameter analysis on $\tau$
@@
Table A6: Parameter analysis results of $\tau$ on model accuracy in the SUTD-TrafficQA dataset.
Appendix A4 Real-world Experiments on the Volleyball Dataset
A4.1 Dataset
The volleyball dateset (Ibrahim et al., 2016) is a video action recognition dataset with 4,830 clips from 55 videos. There are 8 group activity labels, including 4 main activities (set, spike, pass, win-point) that are divided into two subgroups, left and right. Two formats for inputs are provided: RGB videos and keypoints time series. In our setting, we simply use key points as the input. We utilized the ’original’ split of the Volleyball dataset in which all videos were randomly assigned, consisting of 39 training videos and 16 testing videos. We adopt this dataset due to the complex occlusion in the sports which is aligned with our non-invertible generation setting.
A4.2 Implementation Details
The method is implemented using a VAE network. Specifically, the Sequence-to-Step Encoder processes the data by first flattening the features from all time steps. Then, following  (Zhou et al., 2022), we apply a Composer to incorporate the interactions with fine-grained information. Subsequently, we aggregate the contextual information through an MLP, mapping from a space of $\mathbb{R}^{T\times d}$ to $\mathbb{R}^{d}$. The Step-to-Step Decoder is also an MLP network mapping from $\mathbb{R}^{d}$ to $\mathbb{R}^{d}$. We adopt the same Modular Prior network as Table A2. For the implementation of TDRL, the only difference is the removal of temporal dependencies during the encoding process of the model (don’t aggregate the contextual information).
A4.3 Results and Analysis
As shown in Table A7, we observe that CaRiNG achieve consistent performance improvement on both person and group activity accuracy. It indicates that the temporal context is useful in the temporal dynamic modeling. Though the goal of this task is the group activity recognition, we found that the person activity accuracy achieves more improvement. It is not surprising since our method ensures better disentanglement and identification of the latent variables of the group activity, i.e., containing the information of persons.
Method CaRiNG TDRL
Group Activity Top1 Accuracy(%) 93.044 92.895
Group Activity Top3 Accuracy(%) 99.028 98.280
Person Activity Top1 Accuracy(%) 74.551 73.286
Person Activity Top3 Accuracy(%) 98.087 96.634
Table A7: Model accuracy in the Volleyball dataset.
Appendix A5 Related Work
A5.1 Causal Discovery with Latent Variables
@@
A5.2 Nonlinear ICA for Time Series Data
Nonlinear ICA represents an alternative methodology to identify latent causal variables within time series data. Such methods leverage auxiliary data—like class labels and domain indices—and impose independence constraints to facilitate the identifiability of latent variables. To illustrate: Time-contrastive learning (TCL (Hyvarinen & Morioka, 2016)) adopts the independent sources premise and capitalizes on the variability in variance across different data segments. Furthermore, Permutation-based contrastive (PCL (Hyvarinen & Morioka, 2017)) puts forth a learning paradigm that distinguishes genuine independent sources from their permuted counterparts. Furthermore, i-VAE (Khemakhem et al., 2020) utilizes deep neural networks, VAEs, to closely approximate the joint distribution encompassing observed and auxiliary non-stationary regimes. Recent work, exemplified by LEAP (Yao et al., 2022b), has tackled both stationary and non-stationary scenarios in tandem. In the stationary context, LEAP postulates a linear non-Gaussian generative process. For the non-stationary context, it assumes a nonlinear generative process, gaining leverage from auxiliary variables. Advancing beyond LEAP, TDRL (Yao et al., 2022a) initially extends the linear non-Gaussian generative assumption to a nonlinear formulation for stationary scenarios. Subsequently, it broadens the non-stationary framework to accommodate structural shifts, global alterations, and combinations thereof. Additionally, CITRIS (Lippe et al., 2022b, a) champions the use of intervention target data to precisely identify scalar and multi-dimensional latent causal factors. However, a common thread across these methodologies is the presumption of an invertible generative process, a stance that often deviates from the realities of actual data. Besides, (Hartford et al., 2022) demonstrates that under a non-invertible scenario without extra information, identifiability can be only achieved in a subspace where bijective mapping exists. Their work provides additional support for the importance of addressing non-invertibility.
A5.3 Temporal modeling